If 500 mL of a solution containing 1.00 g of barium nitrate (199.33 g/mol) is combined with 500 mL of a solution containing 1.00 g of sodium sulfate (142.03 g/mol), will there be a precipitate? Barium sulfate has a K sp value of 9.1 × 10 −11
Answer #1

(3) JINDAL Date Page : 199:33 w come of Ba NO3) vol. of soln (1) - 0.005_M_ tofel Ud= 500 + 5od = 1L comic of Nag sov 142.03